OpenStudy (anonymous):

A man has some hens & some cows. If the number of heads is 50 & number of feet is 42. The number of cows is ?

OpenStudy (anonymous):

sorry feest are 142 !!

OpenStudy (anonymous):

feets-142

OpenStudy (anonymous):

It's almost 21 hens, remaining are cows

OpenStudy (anonymous):

maybe 11 cows

OpenStudy (angela210793):

a cow has 4 feet and a hen 2 x=no of a cow's feet y=no of a hen's feet ---> we can form a system 4x+2y=149 (1) x+y=50 (2) Tht's wht i think....and i guess u know how to solve this system

OpenStudy (angela210793):

x=cow and y=hen